# Kang Chen♣,♡,∗, Zhihao Liu♢,♡,∗, Tonghe Zhang♯,⊳,∗, Zhen Guo, Si Xu, Hao Lin, Hongzhi Zang, & Xiang Li, Quanlu Zhang, Zhaofei Yu, Guoliang Fan, Tiejun Huang, Yu Wang♠,†, Chao Yu♠,♡,†

### v2

@@ 5.3.1 RL algorithms
(b) Value Loss
(c) Explained Variance
Figure 6: Ablation on the critic structure and placement within Flow-SDE $\pi_{0}$ on the LIBERO-Long, indicating that the critic $V_{\text{vlm}}$ attached after the VLM exhibits superior performance. Furthermore, a four-layer MLP demonstrates stronger regression capability than a one-layer MLP in $V_{\text{expert}}$.
5.3.2 Critic Design
Placement. We compare two critic placement strategies, one positioned after the action expert ($V_{\text{expert}}$) and the other after the VLM ($V_{\text{vlm}}$), with $\pi_{0}$ model on the LIBERO-Long task suite. As illustrated in Fig.˜6, both placements yield comparable performance. However, we observe that $V_{\text{vlm}}$ exhibits slightly superior performance, lower value loss, and higher explained variance, despite not receiving the proprioceptive state as input. This advantage can be attributed to a key difference in their input: $V_{\text{vlm}}$ learns a direct mapping from observation to value, while $V_{\text{expert}}$ must contend with optimization challenges arising from coupled state and noisy action inputs.
Nevertheless, to align with the design of the value function, we maintain the $V_{\text{expert}}$ architecture for the $\pi_{0}$, ensuring that state information is incorporated to calculate the value.
(a) Train
(b) Eval
5.3.3 Stochasticity Injection
Flow-Noise and Flow-SDE provide two distinct approaches for injecting stochasticity. Specifically, Flow-Noise employs a learnable noise network, while Flow-SDE uses a fixed noise level strategy as illustrated in Fig.˜3. To isolate the impact of the injection strategy, we evaluate these two strategies on the LIBERO-Long task suite, with the same Flow-SDE MDP formulation. Since the fixed noise approach does not incorporate an entropy coefficient, we set the entropy bonus for learned noise to 0 to ensure a fair comparison.
We set the fixed noise level to $a=0.5$, and the lower and upper bounds for the learnable noise log-variance to 0.08 and 0.16, respectively. As depicted in Fig.˜7, two noise strategies exhibit similar train and eval performance at step 0, which indicates comparable noise magnitudes. Furthermore, the converged performance affirms the efficiency of both injection methods.
@@ 5.3.3 Stochasticity Injection
(b) Update Time
(c) Explained Variance
Figure 8: Ablation on the MDP formulation within Flow-SDE of $\pi_{0}$ on the LIBERO-Long.
5.3.4 Flow Policy MDP
Flow-Noise and Flow-SDE also differ in their MDP formulation, as shown in Fig.˜2. Built on the standard one-layer MDP, Flow-Noise directly calculates the log-likelihood of the denoised sequence for the policy update. In contrast, Flow-SDE constructs a two-layer MDP by integrating the denoising process with the environment, and further employs a hybrid ODE-SDE sampling technique for acceleration. With the same Flow-SDE noise injection strategy, we evaluate these different frameworks on the LIBERO-Long task suite, as illustrated in Fig.˜8.
While the one-layer formulation converges fastest, all three frameworks achieve similar final performance. In terms of computational cost, the hybrid two-layer paradigm reduces training time by half compared to the standard two-layer approach, thanks to a shorter effective MDP chain that lowers the computational cost per RL update. Moreover, we observe that the one-layer MDP shows no significant speed advantage over the standard two-layer model, as its update stage necessitates re-computing the entire denoising trajectory to calculate the log-likelihood, resulting in comparable computational overhead.
5.3.5 Hyper-parameters
Building on the Flow-SDE with $\pi_{0}$, we investigate the influence of the noise level, denoise step, and action chunk on the LIBERO-Spatial benchmark. We denote the train stage as the phase where the policy generates stochastic actions for exploration, whereas the evaluation stage involves generating deterministic actions. The train and eval success rates for the SFT baseline and the RL fine-tuned model after 100 training steps are presented in Tab.˜6.
Table 6: Ablation study of hyperparameters for Flow-SDE on the LIBERO-Spatial. Performance is reported as task success rate (%). “Train” refers to policy performance during the stochastic rollout phase, whereas “Eval” refers to performance during the deterministic evaluation phase.
Models Stage Hyperparameters
Noise Level Denoise Step Action Chunk
0.2 0.5 0.8 1 2 4 8 5 10 20
SFT Train 62.3 56.0 46.6 9.4 28.3 56.1 62.6 56.0 60.7 70.3
Eval 65.2 65.2 65.2 63.8 64.9 65.2 63.2 65.2 70.5 72.6
RL Train 59.5 93.5 95.3 73.8 90.8 93.5 84.3 93.5 93.3 87.5
Eval 73.1 94.5 98.1 88.5 97.0 94.5 86.7 94.5 95.5 89.2
Noise Level. The noise level $a$ in the Flow-SDE is defined in Eq.˜8, which governs the noise injection magnitude during the denoising process. From Tab.˜6, we observe that the SFT baseline’s eval performance is identical across all noise levels as it relies on deterministic ODE sampling. Its training performance, however, degrades as the noise level increases, which is intuitive as higher noise can disrupt the flow path and lead to an inaccurate marginal action distribution.
Extending this analysis to the RL fine-tuning stage reveals a key trade-off: while lower noise levels mitigate performance degradation induced by policy exploration, the capacity for RL refinement is correspondingly constrained. This trade-off is empirically validated in Fig.˜9, which indicates that training with the minimal noise level $a=0.2$ exhibits instability, manifesting as a significantly higher clip fraction. We attribute this instability to the substantially larger gradient magnitudes induced by the low noise level.
(a) Train
(b) Eval
(c) Clipped Fraction
Denoise Step. The denoise step $K$ defines the number of discretization steps for action generation and is critical for controlling the fidelity of the ODE-to-SDE transition in Eq.˜8. In Tab.˜6, we observe that while all configurations start with similar eval performance, the train success rate plummets at $K=1$, indicating a significant ODE-to-SDE discretization error.
However, as in our noise-level analysis, a larger $K$ is not necessarily optimal. As shown in Fig.˜10, a larger $K$ introduces a clear trade-off: it yields higher rollout performance but complicates the training process due to an increased number of denoising steps.
(a) Train
(b) Eval
@@ 5.3.5 Hyper-parameters
Action chunk. The action chunk refers to the number of consecutive actions the policy executes within a single observation. We ablate the action chunk size across 5, 10, and 20, with results presented in Tab.˜6 and further visualized in Fig.˜11.
While a larger chunk size yields a marginal performance improvement, it also reduces the frequency of policy-environment interactions and hinders accurate reward credit assignment. These factors contribute to less reliable advantage estimation, as reflected in the explained variance metric. Consequently, while a large chunk size may provide a stronger SFT baseline, it ultimately constrains the potential gains from subsequent RL fine-tuning. In conclusion, our analysis reveals a consistent trade-off:
Caveat: Hyperparameters optimized for rollout might induce training instability, impeding potential performance gains from RL.
Therefore, a careful selection of these parameters is essential to achieve a suitable balance between train performance and a stable training process.
5.4 Insights from Large-Scale Training
In this subsection, we elaborate on some empirical insights we gained during RL training.
Hyperparameters. According to the hyperparameters ablation detailed in Sec.˜5.3.5, the performance disparity between the train and eval performance of the initial SFT checkpoint warrants close attention. If this disparity is significant, we recommend either reducing the noise magnitude or increasing the number of denoising steps to mitigate the performance degradation caused by the discrepancy between deterministic and stochastic action generation. Furthermore, as previously established, lower noise levels yield larger gradients, requiring a smaller learning rate to maintain training stability.
@@ 5.4 Insights from Large-Scale Training
Training. In our $\pi_{0.5}$ experiments on the LIBERO-Long benchmark, we observed that the Kullback–Leibler (KL) divergence metric increased steadily throughout training, potentially leading to instability. We mitigated this issue by implementing a learning rate scheduler with cosine annealing. As demonstrated in Fig.˜12, this scheduler effectively prevents the KL divergence from escalating, thereby stabilizing the training process.
(a) Eval
(b) KL Divergence
Figure 12: Ablation study on the learning rate scheduler. The experiment is conducted with Flow-SDE $\pi_{0.5}$ on the LIBERO-Long benchmark, demonstrating that the scheduler alleviates over-optimization and stabilizes the training process.
@@
Critic. In our ManiSkill experiments, we observe that policy evaluation performance exhibits an initial dip before improving for both $\pi_{0}$ and $\pi_{0.5}$ models, as shown in Fig.˜14. We attribute this transient degradation to the critic providing inaccurate signals during its warm-up phase. The subsequent eval improvement correlates directly with the critic’s value estimations stabilizing, as evidenced by the rising explained variance.
@@
Temporal Efficiency We also study how the rollout of RL in a physical simulator helps shape the policy to achieve expert-level temporal efficiency. We analyze the expert motion planning data used for SFT and then tracked the average episodes lengths during the RL training of the $\pi_{0.5}$ model using our methods. As shown in Figure Fig.˜13, the SFT-initialized policy exhibits significantly longer episodes due to execution errors. In contrast, $\pi_{0.5}$ achieves episode lengths that converge to the expert range after RL training, demonstrating a substantial improvement in temporal efficiency.
We attribute this convergence to two factors: (1) RL’s error-correction capability, which helps the policy succeed in a more diverse distribution, and (2) our partial reset mechanism with discounted reward, where faster task completion leads to more resets and higher cumulative reward between updates.
Figure 13: Episode length: $\pi_{0.5}$ RL training in multi-object pick-and-place environment
(a) Eval
(b) Explained Variance
Figure 14: Flow-Noise Training curve in ManiSkill generalization test.
5.5 Extension: Fine-tune VLM and Action Expert Simultaneously
In our previous experiments, the VLM is frozen, and the optimization is confined exclusively to the action expert during RL. In this subsection, we aim to investigate the role of the VLM during RL. Specifically, we employ Low-Rank Adaptation (LoRA) (hu2022lora) for the VLM, facilitating its joint optimization with the action expert. We set the LoRA rank to $r=32$ and the scaling parameter to $\alpha=32$, while the action expert remains fully trainable.
We conduct experiments with the $\pi_{0}$ model with Flow-SDE on the LIBERO-Long benchmark, comparing three distinct configurations: 1) VLM frozen baseline ($5\times 10^{-6}$ learning rate, 4 updates per epoch), 2) VLM LoRA-I ($5\times 10^{-6}$ learning rate, 4 updates per epoch), and 3) VLM LoRA-II with conservative update training config ($1\times 10^{-6}$ learning rate, 2 updates per epoch). As presented in Fig.˜15, the VLM LoRA-II configuration achieves a learning trajectory comparable to the VLM frozen baseline. This empirical observation yields two critical inferences: First, the benefit of fine-tuning the VLM on the LIBERO benchmark is not evident; Second, fine-tuning VLM together with the action expert requires a more conservative optimization configuration for training stability. We conjecture the limited performance gain attributable to the limited scene variability within LIBERO, for which the pretrained VLM representations are already sufficiently robust.
(a) Eval
(b) KL Divergence
Figure 15: Ablation study on VLM Effectiveness during RL. The experiment is conducted with Flow-SDE $\pi_{0}$ on the LIBERO-Long benchmark. We compare the performance of a frozen VLM baseline (learning rate $5\times 10^{-6}$, 4 updates per epoch) against two LoRA-tuned VLM configurations: LoRA-I (using the same training config) and LoRA-II (a more conservative setting with learning rate $1\times 10^{-6}$ and 2 updates per epoch).
6 Conclusion
We introduce $\pi_{\texttt{RL}}$, the first framework that enables flow-based VLAs, $\pi_{0}$ and $\pi_{0.5}$, to be fine-tuned with PPO. We tackle the fundamental challenge of intractable log-likelihoods in flow matching and propose two technical solutions, Flow-Noise and Flow-SDE, which differ in their stochasticity injection strategies and MDP formulations. Our extensive experiments on the challenging LIBERO and ManiSkill benchmarks demonstrated that $\pi_{\texttt{RL}}$ achieves significant performance improvements over SFT baselines.
7 Limitations and Future Work
@@
Noise Injection. Our current noise injection strategy exhibits some train performance drop during the ODE-to-SDE conversion. Flow-CPS (wang2025coefficients) attributes this loss to numerical error and proposes an improved coefficients-preserving sampling method. In our experiments, we attempted this configuration. Consistent with our hyperparameter ablation, our experiments showed that while this configuration mitigated the ODE-SDE precision error, it yielded limited RL improvement. Nevertheless, we argue that improving the noise injection strategy holds significant potential, specifically converting the ODE formulation to an SDE formulation while preserving the action distribution undisturbed.
@@
Training Acceleration. Our current implementation of the mixed ODE-SDE rollout is simplistic in Flow-SDE, i.e., it randomly selects one denoising step as an SDE step, while all other steps remain ODE steps. We posit that future investigations into mixed ODE-SDE rollouts, leveraging advances in accelerating flow-based image generation (li2025mixgrpo; he2025tempflow; liu2025flowgrpo; li2025branchgrpo), could further enhance Flow-SDE, leading to faster training and improved performance.
Generalization. Our experiments in the Maniskill OOD tests indicate that the semantic generalization capabilities of the SFT and RL models remain limited. We aim to investigate and improve this issue in future studies.
Real-world Experiment. Our current experiments are evaluated solely in simulated environments, lacking empirical validation in a physical system. We plan to extend this research by applying our RL methodology to real-world tasks in the future.
Appendix A Algorithm Details
GRPO is a critic-free method that estimates the advantage by normalizing rewards within a group of rollouts from the same state. In our robotics MDP task, for each initial state, we use the policy $\pi_{\theta}$ to sample a group of $G$ trajectories, resulting in $G$ sparse terminal rewards $\{\mathcal{R}^{(j)}\}_{j=1}^{G}$ denoting the binary success of the task. The advantage for the $i$-th trajectory, $\hat{A}^{(i)}$, is then calculated based on the group-wise reward normalization:
$\hat{A}^{(i)}=\frac{\mathcal{R}^{(i)}-\text{mean}(\{\mathcal{R}^{(j)}\}_{j=1}^{G})}{\text{std}(\{\mathcal{R}^{(j)}\}_{j=1}^{G})}$ (17)
where $\mathcal{R}^{(i)}$ is the terminal reward for the $i$-th trajectory, and the mean and standard deviation are computed over the group of $G$ trajectories. Since the reward is only granted at the end of an episode, the advantage estimate remains constant across all timesteps within that trajectory.
Appendix B Experiment Details
@@
We record the training hyperparameters used to train both $\pi_{0}$ and $\pi_{0.5}$ on each LIBERO task, and present them in Tabs.˜7 and 8.
Table 7: Hyperparameters of Flow-Noise and Flow-SDE with PPO across LIBERO tasks.
@@
Train epochs 400 400 400 400 400 400 400 400
Batch size 2048 2048 2048 2048 2048 2048 2048 2048
Update epochs 2 2 4 4 1 1 3 4
Actor lr 1e-5 5e-6 5e-6 5e-6 5e-6 5e-6 5e-6 5e-6
@@
Action chunk $H$ 5 5 5 10 5 5 5 10
@@
Table 8: Flow-Noise and Flow-SDE Hyperparameters of in ManiSkill tasks.
@@
Eggplant Carrot Spoon Cube Generalization Eggplant Carrot Spoon Cube Generalization
SFT train steps 1000 100040 1000 1000 1000 1000 1000 1000 1000 1000
@@
w
Table 9: $\pi_{0}$, $\pi_{0.5}$ Generalization Test
Flow-SDE Flow-Noise Flow-SDE Flow-Noise
In distribution Main-v3-train 38.42 78.83 77.76 40.06 90.85 89.65
Visual-language
Variations
Instruct-v1-test 30.10 64.58 66.46 46.56 76.98 85.73
VisionImage-v1-test 38.33 68.75 71.67 46.25 78.75 83.13
VisionTexture03-v1-test 35.10 66.04 66.77 36.67 69.58 75.00
VisionTexture05-v1-test 31.04 55.83 60.52 32.71 58.02 62.19
VisionWhole03-v1-test 35.42 62.40 68.96 40.10 69.58 71.56
VisionWhole05-v1-test 28.54 48.96 53.85 30.73 55.00 56.98
Semantic Reasoning
(object/receptacle
confounders)
MultiCarrot-v1-test 7.81 28.23 23.02 16.67 36.77 38.23
MultiCarrot-v1-train 12.50 36.46 31.77 28.23 49.48 50.10
MultiPlate-v1-test 5.00 16.35 18.33 11.77 29.38 28.33
MultiPlate-v1-train 7.29 20.52 19.58 9.69 22.29 25.42
Action Execution PositionChangeTo-v1-test 9.58 17.40 10.94 13.54 36.25 54.69
Position-v1-test 16.88 45.63 37.50 31.15 54.48 55.00
Appendix C Additional Results: RL for GR00T N1.5
C.1 Setup
GR00T N1.5. We conduct additional experiments based on the GR00T N1.5 model. GR00T N1.5 is an open-source foundation model  [^8] https://github.com/NVIDIA/Isaac-GR00T tailored for generalist humanoid robot reasoning and manipulation. Designed to enable cross-embodiment adaptability, the model accepts multimodal inputs—including natural language instructions and visual observations—and integrates robot state information (e.g., joint positions, end-effector poses) to generate continuous motor actions for diverse tasks and environments. Its neural architecture combines a vision-language model (Eagle 2.5) optimized for grounding and physical understanding with a diffusion transformer (DiT) head for action denoising. It supports multiple robot embodiments via specialized heads—including humanoid robots with dexterous hands, single-arm robots, and humanoid robots with grippers. For the critic placement, we estimate the value across the entire denoising trajectory, attaching the critic network to the action head. The figure illustration is shown in Fig.˜16.
Benchmark. We use LIBERO as the evaluation benchmark. Evaluates the model performance across four manipulation task suites: Spatial, Object, Goal, and Long.
Implementation Details. Similar to the $\pi_{0}$ implementation, we initiate our process with SFT on expert demonstrations. For the SFT stage, we fine-tune the entire model following the official setting. In the subsequent RL stage, we exclusively fine-tune the action expert model while keeping the vision-language model parameters fixed. A critical methodological detail during the RL stage is the replacement of the original dropout layers in the expert model with identity layers. Dropout layers are known to destabilize online RL training. Specifically, they introduce an unintended change in the effective policy, transforming the stable probability ratio update from:ρ_t(θ) = πθnew(at|st)πθold(at|st)to the unstable form:ρ_t(θ) = παnew(at|st)πθold(at|st),where $\rho_{t}(\theta)$ is the probability ratio, and $\alpha_{\text{new}}$ signifies the policy after the update and the non-deterministic effect of the dropout mask. This stochasticity, which effectively changes the model’s structure in addition to the per-step policy update, significantly deteriorates training stability. For the model configurations, we adhere to the official settings. Our experiments are conducted on 8 NVIDIA H100 80GB GPUs, and detailed training hyperparameters are exactly the same as $\pi_{0}$, listed in Tab.˜7. The RL training method selected is Flow-SDE.
Figure 16: Illustration for the architecture of GR00T-N1.5.
Table 10: Comparison of the PPO and GRPO with Flow-SDE on the LIBERO.
Model LIBERO
Spatial Object Goal Long Avg. $\Delta$ Avg.
$GR00T$ SFT 41.4 58.6 48.2 61.9 52.5
+PPO 92.5 96.2 84.3 86.6 89.9 +37.4
C.2 Results
This evaluation adopts all the hyperparameter settings of $\pi_{0}$ without further tuning to show the broad applicability and robustness of the proposed RL training method. The evaluation results are shown in Tab.˜10 For the few-shot model, the SFT baseline performs poorly, with an average success rate of only 52.5%, indicating that the model struggles with limited demonstration data. Our proposed RL training method substantially boosts performance, reaching 89.9% with Flow-SDE.
The results presented above utilize the identical hyperparameter settings as $\pi_{0}$. These findings primarily serve to demonstrate the broad applicability and inherent robustness of the proposed RL training methodology. Further optimization through parameter tuning is likely to yield enhanced model performance.

### v3

@@ E.1 SIMPLER
(b) KL Divergence
Figure 13: Ablation study on the effectiveness of VLM during RL.
@@
(b) KL Divergence
Figure 14: Ablation study on the learning rate scheduler.
E.2 Real2Sim2Real
While the SIMPLER benchmark demonstrates predictive correlation between simulation and real-world performance, a pronounced visual domain shift remains to be solved. To this end, we leverage recent Real2Sim2Real methodologies (Fan et al., 2025; Jiang et al., 2025) to construct a high-fidelity simulation environment with ManiSkill for rigid body dynamics and Gaussian Splatting (Kerbl et al., 2023) for photorealistic rendering.
Setup. Our hardware platform comprises a Franka Panda robotic arm and an Intel RealSense D435 camera serving as the primary visual sensor. We perform manual calibration by aligning simulated viewpoints with real-world camera perspectives to synchronize their extrinsic matrices. As illustrated in Fig.˜1, the photorealistic textures and color profiles in our simulator closely mirror the physical environment, effectively minimizing the visual domain shift from simulation to reality.
Results. Following the experimental protocol established in the ManiSkill benchmarks, we initially collect 20 expert trajectories via a motion planner for few-shot SFT, which is subsequently optimized through RL over 100 training iterations. We deploy the RL fine-tuned policies in the real world in a zero-shot manner. Empirical results indicate that while the SFT baseline fails to complete the task, the RL-tuned policy achieves a 40% success rate. A representative successful episode is visualized in Fig.˜12.
Appendix F Ablation Details
F.1 RL algorithms
Given the significant performance gains from PPO on the LIBERO benchmark, we also investigated the effectiveness of GRPO (Shao et al., 2024), another widely used policy gradient method applied in VLA+RL training (Li et al., 2025a). We compare the performance of PPO and GRPO on both the $\pi_{0}$ and $\pi_{0.5}$ models, with results denoted in Tab.˜9.
Conclusion. To sum up, our findings highlight a critical trade-off: parameters tailored for rollout success may adversely impact training stability, ultimately constraining the performance ceiling of RL. Therefore, careful parameter tuning is required to achieve a synergy between high-quality rollouts and stable policy convergence.
@@ F.1 RL algorithms
Figure 15: Training curves in ManiSkill.
Figure 16: Episode length: $\pi_{0.5}$ RL training in ManiSkill.
F.2 VLM Fine-tuning Analysis
In our previous experiments, the VLM is frozen, and the optimization is confined exclusively to the action expert during RL. In this subsection, we aim to investigate the role of the VLM during RL. Specifically, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) for the VLM, facilitating its joint optimization with the action expert. We set the LoRA rank to $r=32$ and the scaling parameter to $\alpha=32$, while the action expert remains fully trainable.
We conduct experiments with the $\pi_{0}$ model with Flow-SDE on the LIBERO-Long benchmark, comparing three distinct configurations:
  • VLM Frozen: $5e^{-6}$ learning rate, 4 updates/epoch.
  • VLM LoRA-I: $5e^{-6}$ learning rate, 4 updates/epoch.
  • VLM LoRA-II: $1e^{-6}$ learning rate, 2 updates/epoch.
As presented in Fig.˜13, the VLM LoRA-II configuration achieves a learning trajectory comparable to the VLM frozen baseline. This empirical observation yields two critical inferences: (1) The benefit of fine-tuning the VLM on the LIBERO benchmark is not evident. We conjecture the limited performance gain owing to the limited scene variability within LIBERO, for which the pretrained VLM representations are already sufficiently robust. (2) Fine-tuning VLM together with the action expert requires a more conservative optimization configuration for training stability.
Appendix G Insights from Large-Scale RL Training
In this section, we elaborate on some empirical insights we gained during RL training.
Hyperparameters. According to the hyperparameters ablation detailed in Sec.˜5.4, the performance disparity between the train and eval performance of the initial SFT checkpoint warrants close attention. If this disparity is significant, we recommend either reducing the noise magnitude or increasing the number of denoising steps to mitigate the performance loss when shifting from deterministic to stochastic execution. Furthermore, as previously established, lower noise levels yield larger gradients, requiring a smaller learning rate to maintain training stability.
@@ Appendix G Insights from Large-Scale RL Training
Training. In our $\pi_{0.5}$ experiments on the LIBERO-Long benchmark, we observed that the Kullback–Leibler (KL) divergence metric increased steadily throughout training, potentially leading to instability. We mitigated this issue by implementing a learning rate scheduler with cosine annealing. As demonstrated in Fig.˜14, this scheduler effectively prevents the KL divergence from escalating, thereby stabilizing the training process.
@@
Critic. In our ManiSkill experiments, we observe that policy evaluation performance exhibits an initial dip before improving for both $\pi_{0}$ and $\pi_{0.5}$ models, as shown in Fig.˜15. We attribute this transient degradation to the critic providing inaccurate signals during its warm-up phase. The subsequent eval improvement correlates directly with the critic’s value estimations stabilizing, as evidenced by the rising explained variance.
@@
Temporal Efficiency. We also study how the rollout of RL in a physical simulator helps shape the policy to achieve expert-level temporal efficiency. We analyze the expert motion planning data used for SFT and then tracked the average episode lengths during the RL training of the $\pi_{0.5}$ model in ManiSkill. As shown in Fig.˜16, the SFT-initialized policy exhibits significantly longer episodes due to execution errors. In contrast, $\pi_{0.5}$ achieves episode lengths that converge to the expert range after RL training, demonstrating a substantial improvement in temporal efficiency.
We attribute this convergence to two factors: (1) RL enhances the policy’s error-correction capabilities, allowing it to recover from execution failures. (2) Our partial reset mechanism incentivizes temporal efficiency through discounted rewards, as faster task completion enables the agent to trigger more resets and accumulate higher total rewards within each update cycle.
Figure 17: Illustration for the architecture of GR00T-N1.5.
Table 10: Results of Finetuning GR00T using PPO with Flow-SDE on the LIBERO.
Model LIBERO
Spatial Object Goal Long Avg. $\Delta$ Avg.
GR00T SFT 41.4 58.6 48.2 61.9 52.5
+PPO 92.5 96.2 84.3 86.6 89.9 +37.4
Appendix H RL for GR00T N1.5
H.1 Setup
GR00T N1.5. We conduct additional experiments based on the GR00T N1.5 model (Bjorck et al., 2025), which is a foundation model tailored for generalist humanoid robot reasoning and manipulation. The architecture integrates an Eagle 2.5 VLM (Chen et al., 2025), optimized for spatial grounding and physical reasoning, with a Diffusion Transformer head (Peebles and Xie, 2023) for action denoising. It facilitates multi-embodiment compatibility through specialized heads, supporting configurations such as humanoids with dexterous hands or grippers, as well as single-arm manipulators.
Regarding the critic implementation, we estimate value functions across the entire denoising trajectory by integrating the critic network directly with the action head. The complete framework is illustrated in Fig.˜17.
Benchmark. We evaluate the model performance of GR00T across four manipulation task suites in LIBERO: Spatial, Object, Goal, and Long.
Implementation Details. Similar to the $\pi_{0}$ implementation, we initiate our process with SFT on expert demonstrations. For the SFT stage, we fine-tune the entire model following the official setting. In the subsequent RL stage, we exclusively fine-tune the action expert model while keeping the vision-language model parameters fixed.
A crucial methodological refinement in our RL pipeline is the replacement of dropout layers in the expert model with identity layers. Dropout is widely recognized to induce instability during online RL training. Specifically, it introduces non-deterministic perturbations to the effective policy, shifting the standard probability ratio update from:
$\rho_{t}(\theta)=\frac{\pi_{\theta_{\text{new}}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ (17)
to a highly unstable form:
$\rho_{t}(\theta)=\frac{\pi_{\alpha_{\text{new}}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},$ (18)
where $\rho_{t}(\theta)$ denotes the probability ratio, and $\alpha_{\text{new}}$ represents the policy state post-update as modified by the stochastic dropout mask. This structural stochasticity, compounded with per-step policy updates, severely undermines training convergence. The training hyperparameters are identical to those used for $\pi_{0}$, with Flow-SDE employed as the primary RL algorithm.
H.2 Results
Results are summarized in Tab.˜10. For the few-shot model, the SFT baseline achieves only a 52.5% success rate, reflecting limited generalization from sparse demonstrations. Conversely, our RL-based Flow-SDE significantly improves performance to 89.9%. These results, obtained using default $\pi_{0}$ configurations, underscore the broad applicability of our method across architectures. While task-specific tuning could further enhance performance, we leave such optimization for future work.
Appendix I Limitations and Future Work
@@
Noise Injection. Our current noise injection strategy exhibits a performance drop during the ODE-to-SDE conversion. Flow-CPS (Wang and Yu, 2025) attributes this loss to numerical error and proposes an improved coefficients-preserving sampling method. In our experiments, we attempted this configuration. Consistent with our hyperparameter ablation, our experiments showed that while this configuration mitigated the ODE-SDE precision error, it yielded limited RL improvement. Nevertheless, we argue that improving the noise injection strategy holds significant potential, specifically converting the ODE formulation to an SDE formulation while preserving the action distribution undisturbed.
@@
Training Acceleration. Our current implementation of the mixed ODE-SDE rollout is simplistic in Flow-SDE, i.e., it randomly selects one denoising step as an SDE step, while all other steps remain ODE steps. We posit that future investigations into mixed ODE-SDE rollouts, leveraging advances in accelerating flow-based image generation (Li et al., 2025b; He et al., 2025; Liu et al., 2025a; Li et al., 2025c), could further enhance Flow-SDE, leading to faster training and improved performance.
Generalization. Maniskill OOD tests indicate that the semantic generalization of SFT and RL models remains limited. To address this, future work will leverage RL to enhance robustness by training on more diverse task distributions and varied linguistic instructions, thereby fostering better cross-task adaptability.
Appendix J Training Hyperparameters.
@@
We record the training hyperparameters used to train both $\pi_{0}$ and $\pi_{0.5}$ on each benchmark, and present them in Tabs.˜11, 12 and 13.
Table 11: Hyperparameters across LIBERO.
@@
Train epochs 500 500 500 500 500 500 500 500
Global batch size 2048 2048 2048 2048 2048 2048 2048 2048
Update epochs 4 4 4 4 1 1 3 4
Actor lr 5e-6 5e-6 5e-6 5e-6 5e-6 5e-6 5e-6 5e-6
@@
Action prediction horizon $H$ 50 50 50 50 10 10 10 10
Action replan horizon $H^{\prime}$ 5 5 5 10 5 5 5 10
@@
Table 12: Hyperparameters across SIMPLER and ManiSkill.
@@
Eggplant Carrot Spoon Cube ManiSkill Eggplant Carrot Spoon Cube ManiSkill
SFT train steps 1000 1000 1000 1000 1000 1000 1000 1000 1000 1000
@@
Table 13: Hyperparameters across MetaWorld and CALVIN benchmarks.
Parameters Benchmarks and models
MetaWorld CALVIN
$\bm{\pi_{0}}$ $\bm{\pi_{0.5}}$ $\bm{\pi_{0}}$ $\bm{\pi_{0.5}}$
Train epochs 450 450 100 100
Global batch size 2048 2048 2048 2048
Update epochs 4 4 4 4
Actor lr 1e-5 5e-6 5e-6 5e-6
Critic lr 1e-4 1e-4 1e-4 1e-4
Scheduler False True False False
Reward discount rate $\gamma$ 0.99 0.99 0.99 0.99
GAE $\lambda$ 0.95 0.95 0.95 0.95
Clip ratio $\epsilon$ 0.2 0.2 0.2 0.2
Interaction steps 100 100 480 480
Parallel environments 64 64 64 64
Rollout epochs 8 8 8 8
Action prediction horizon $H$ 5 5 5 5
Action replan horizon $H^{\prime}$ 5 5 5 5
Denoise steps 5 5 5 5
Noise level $\sigma$ (Flow-SDE) 0.5 0.5 0.5 0.5
Max log-var (Flow-Noise) 0.10 0.10 0.16 0.16
Min log-var (Flow-Noise) 0.04 0.04 0.08 0.08
Entropy bonus (Flow-Noise) 0.005 0.005 0.005 0.005